OpenStudy (anonymous):

Select the expression equal to (radical sign) 75 A. 5 (R sign) 7 B. 7 (R sign) 5 C. 5 (sign) 3 D. 3 (sign) 5. (How do i solve this)?

OpenStudy (kenljw):

(75^1/2 = (3X25)^1/2 = 5(3^1/2)

OpenStudy (anonymous):

Can you explain the last part Kenl)w

OpenStudy (kenljw):

radical can be expressed as fraction exponents 75 = 25X3 75^1/2 = (25^1/2)(3^/1/2) = 5 (3^1/2)

OpenStudy (anonymous):

Oh, Okay what would the answer be?

OpenStudy (kenljw):

(ab)^n = a^n b^n

OpenStudy (kenljw):

C

OpenStudy (anonymous):

Thank you for your help (: